### v1

# Real Time Proportional Throughput Maximization: How much advance notice should you give your scheduler?

Real Time Proportional Throughput Maximization: How much advance notice should you give your scheduler?
Nadim A. Mottu
(November 20, 2025)
1 Introduction
1.1 Abstract
We will be exploring a generalization of real time scheduling problem sometimes called the real time throughput maximization problem. Our input is a sequence of jobs specified by their release time, deadline and processing time. We assume that jobs are announced before or at their release time. At each time step, the algorithm must decide whether to schedule a job based on the information so far. The goal is to maximize the value of the sum of the processing times of jobs that finish before their deadline, or the total “active” time.
We extend this problem by defining a notion of $t$-advance-notice, a measure of how far in advance each job is given relative to their processing time. We show that there exists a $\frac{t}{2t+1}$-competitive algorithm when all jobs have $t$-advance-notice for $t\in[0,1]$. We also show that this ratio is optimal for all algorithms with $t$-advance-notice and that the upper bound of $\frac{t}{2t+1}$-competitiveness holds for all $t$, in particular that regardless of how much advance-notice is given, no algorithm can reach $\frac{1}{2}$-competitiveness.
1.2 Our model
Our input is a sequence of jobs $\mathcal{J}=(J_{1},\ldots,J_{n})$ where $J_{i}=(a_{i},r_{i},p_{i},d_{i})\in\mathbb{R}_{\geq 0}^{4}$. In this model, each job is announced by an adversary at time $a_{i}=a(J_{i})$, at this time, its “release time” $r_{i}=r(J_{i})$, processing time $p_{i}=p(J_{i})$ and deadline $d_{i}=d(J_{i})$ are revealed to the algorithm. The algorithm does not know what jobs may be announced in the future or the number of jobs $n$. At each time step, the algorithm ALG must decide whether to schedule a job based on the information so far, subject to the restriction that if a job $J_{i}$ is scheduled at time $s_{i}=s(J_{i})=s^{\textsc{ALG}}(J_{i})$ then no other job can be scheduled during the interval $s_{i},s_{i}+p_{i})$, furthermore, $r_{i}\leq s_{i}\leq s_{i}+p_{i}\leq d_{i}$. The goal is to maximize the value of the sum of the processing times $p_{i}$ of jobs that finish before their deadline, or the total “active” time. Later in this paper, we will also briefly explore the alternative goal of maximizing the total number of jobs that finish before their deadline (regardless of length), we call this case “unweighted”. Both these objectives are special cases of a “weight” maximization where we aim to maximize the total “weight” of jobs that finish before the deadline, in the former case, we assume that the weight is proportional to the processing time $p_{i}$ of each job, and in the latter we assume the weight is some fixed constant. We will also briefly look at maximizing the value of other weight functions such as $C$-benevolent (convex increasing) functions.
1.3 Related Works
The problem of real time throughput and it’s applications has been studied quite thoroughly.
It is not difficult to show that in general, any algorithm can be made to perform arbitrarily bad. After all, an adversarial scheduler could schedule a job of weight 1, then schedule another job of arbitrarily large weight overlapping with that job if and only if the algorithm takes the original job. Regardless of what the algorithm does, its performance when compared to an optimal offline algorithm is arbitrarily large. For this reason, most of the existing work related to this problem tackle special cases or restrictions on the problem.
The problem of job scheduling with fixed start and end times has been studied by Woeginger [WOEGINGER19945] where he found that maximizing the total weights of jobs could be done by a $\frac{1}{4}$-competitive algorithm when the weights are assumed to be given by functions he describes as $C$-benevolent and $D$-benevolent and the algorithm is allowed to “revoke” another job, stopping it’s execution early and forfeiting it’s value in exchange for scheduling another, better job. Notably, since the identity function (and any linear function) is $C$-benevolent, his result applies to the proportional case as well. The case where start and end times are fixed is often referred to as “interval scheduling”, as opposed to “job scheduling”. Later, still focusing on intervals Fung et al. [Fung2014] gives a $\frac{1}{2}$-competitive ratio barely random algorithm for “monotone” intervals as well as $C$-benevolent and $D$-benevolent weight functions. He also give a $\frac{1}{3}$-competitive algorithm for equal length jobs allowing for variable start times (also known as slack). In [Fung2014], he assumes that algorithms have the ability to “restart” other jobs, which is to say interrupt them similar to “revoking”, only now we also allow jobs that are revoked to be started again from scratch so long as their deadline has not past. This work built on previous results from Chrobak et al. [ChrobakRestarts] who showed that for equal processing time jobs deterministic algorithms a competitive ratio of $\frac{2}{3}$ is tight when restarting is allowed for unweighted job scheduling. However, they also show that when inputs are given sufficiently large “slack”, restarting is not needed for their algorithm to reach the desired competitive ratio. The more general case of restarting algorithms for unweighted job scheduling is studied by Hoogeveen et al. [HOOGEVEEN2000193] where he gives a tight $\frac{1}{2}$ competitive ratio for arbitrarily processing times in the unweighted restart model.
On the offline side of the problem, [BaptisteOn4] Baptiste gives a polynomial time and space algorithm for maximizing the number of on time jobs. As for the weighted case, the very closely related task of minimizing weight of late jobs is shown to be NP-Complete by Karp [Karp1972]. A $(1-\frac{1}{e}-\varepsilon)$-approximation is given by Karakostas [KARAKOSTAS2025115258]. Meanwhile, Borodin et al. [Borodin2003] show that there exists a priority algorithm with $\frac{1}{3}$-competitive ratio for unweighted job scheduling and $\frac{1}{4}$-competitive ratio for proportional weights.
In general, when studying real time scheduling, the time of announcement $a_{i}$ and the time of release $r_{i}$ are assumed to be the same for all jobs. That is to say, to the best of our knowledge, we are the first to explore a model where jobs may be announced ahead of their release time in a real-time model. The main benefit of this addition is that it allows us to reach a finite, and relatively good, competitive ratio without the need to restart or preempt. The problem of throughput in a real time context is very well motivated by many applications such as CPU scheduling,among other things. In many contexts, however, it may be inappropriate or detrimental to stop or revoke a job in the middle of it’s execution. Somebody who booked a room, or a service, may for example find it inconvenient to lose access in the middle of their alloted time. It may, in these circumstances, seem far more reasonable to require users or “jobs” to announce their availability and required time in advance. Ultimately, this extension feels very natural, and provides what we view as an appropriate alternative to the preemption models studied elsewhere.
1.4 Preliminary
Our goal will be to design algorithms to maximize an (asymptotic) competitive ratio defined as follows:
$\rho(\textsc{ALG})=\lim_{\textsc{OPT}(\mathcal{J})\to\infty}\inf\frac{\textsc{ALG}(\mathcal{J})}{\textsc{OPT}(\mathcal{J})}$
Where $\textsc{OPT}(\mathcal{J})$ denotes the value of the optimal offline solution on a sequence of jobs $\mathcal{J}=(J_{1},\ldots,J_{n})$, and $\textsc{ALG}(\mathcal{J})$ is the value of the algorithm on that same set of jobs.
It is clear that in general, the inclusion of an announcement time does not help us. Since giving advance notice $a_{i}<r_{i}$ does not impact the value of the optimal offline solution (which knows about every job from the very beginning), an adversary designing the sequence of jobs $\mathcal{J}$ can always make an algorithm perform worst by making $a_{i}=r_{i}$ without impacting the optimal solution. We therefore need to place a restriction on the value of $a_{i}$ to allow our algorithms to be able to reliably use it.
Definition 1.1.
We say that a job $J$ has $t$-advance-notice for some constant $t\in\mathbb{R}_{\geq 0}$ if $(r_{i}-a_{i})\geq t\cdot p_{i}$.
Since jobs are designed by an adversary and giving an earlier $a_{i}$ only benefits an online algorithm without changing the optimum, we can assume in most instances that $(r_{i}-a_{i})=t\cdot p_{i}$ (the adversary does not give any more early notice than what is strictly required).
2 The power and limitations of advance-notice
In this section, we will provide a construction for an algorithm that achieves $\frac{t}{2t+1}$-competitive ratio on job sequences with $t$-advance-notice for $t\in[0,1]$. We also show that this algorithm is optimal by showing that any algorithm on a sequence of jobs with $t$-advance-notice that does not use revoking must be at most $\frac{t}{2t+1}$-competitive for proportional weights without preemption. As a consequence, we show that for any $c\in[0,\frac{1}{3}]$, there exists a $t$ such that we can achieve $c$-competitiveness when we are guaranteed $t$-advance-notice but that for any $c\in[\frac{1}{2},1]$, we can never achieve $c$-competitiveness regardless of advance-notice.
2.1 The power of advance-notice
We begin by showing that we can achieve $\frac{t}{2t+1}$-competitiveness when jobs have $t$-advance-notice. We do this by constructing an offline algorithm Off and running it on all the information received so far whenever a new job is announced.
Theorem 2.1.
Let Off be any optimal offline algorithm for proportional throughput. There exists an online algorithm $\mathcal{A}_{\textsc{Off}}$ which is at least $\frac{t}{2t+1}$-competitive for real time proportional throughput when all jobs have $t$-advance-notice for $t\in[0,1]$.
Proof:
[…]
Algorithm 1 code $\mathcal{A}_{\textsc{Off}}$
[…]
4:for timestep $t$ do
5:  if $(t,J)\in S$ for some $J\in\mathcal{J}$ then
6:   $s(J)\leftarrow t$ $\triangleright$ Schedule $J$
7:   $J^{*}\leftarrow J$
8:  if $t+p(J)>d(J)$ for some $J\in\mathcal{J}$ then $\triangleright$ $J$ is no longer feasible
9:   $\mathcal{J}\leftarrow\mathcal{J}\setminus\{J\}$
10:  if $s(J^{*})+p(J^{*})\geq t$ then $\triangleright$ Current running job is done
11:   $J^{*}\leftarrow\bot$
12:  if $J$ is announced at time $t$ then
13:   $L\leftarrow\max\{s(J^{*})+p(J^{*}),t\}$ $\triangleright$ The latest time we can schedule a new job
14:   $\mathcal{J}^{\prime}\leftarrow\{J=(a(J),\max\{r(J),L\},p(J),d(J)):J\in\mathcal{J}\}$ $\triangleright$ feasible jobs with start cut off
15:   $S\leftarrow\textsc{Off}(\mathcal{J}^{\prime})$
In other words, at each announcement, run Off on the set of known jobs not run yet, possibly cutting off their release time to only the feasible times.
We will show that this algorithm is $\frac{t}{2t+1}$-competitive by a charging argument. Let $\textsc{OPT}(\mathcal{J})$ be the set of jobs scheduled by OPT when run on the whole set of jobs and let $\mathcal{A}_{\textsc{Off}}(\mathcal{J})$ be the set of jobs run by $\mathcal{A}_{\textsc{Off}}$.
Let $s_{i}$ and $s(J_{i})$ be the time a job $i$ is scheduled by $\mathcal{A}_{\textsc{Off}}$ and let $s_{i}^{\textsc{OPT}}$ along with $s^{\textsc{OPT}}(J_{i})$ be the time job $i$ is scheduled in the optimal solution OPT
We say that a job $J_{i}$ in OPT “conflicts” with a job $J_{j}$ in $\mathcal{A}_{\textsc{Off}}$ if $s_{j},s_{j}+p_{j})\cap[s_{i}^{\textsc{OPT}},s_{i}^{\textsc{OPT}}+p_{i})\neq\emptyset$.
We will use the following charging argument:
For each job in $J$ in $\textsc{OPT}(\mathcal{J})$ we will assign them to the following ways:
  • If it does not conflict with a job in $\mathcal{A}_{\textsc{Off}}(\mathcal{J})$ then we assign it to itself. In such a case we know that $\mathcal{A}_{\textsc{Off}}$ would have scheduled $J$ at a different point, otherwise it would have scheduled $J$ at the same point at OPT seeing as it is feasible at that point and available. We call the set of jobs assigned to $J_{i}$ by this way $A_{i}$.
  • If it conflicts with a job $J_{i}$ in $\mathcal{A}_{\textsc{Off}}$ and is contained within $J_{i}$,
  • If it conflicts with a job $J_{i}$ in $\mathcal{A}_{\textsc{Off}}$ announced during $J_{i}$, then it is assigned to that job $J_{i}$. The set of all such jobs is $C_{i}$
  • If it conflicts with only 1 job in $\mathcal{A}_{\textsc{Off}}$ $J_{i}$ then it is assigned to $J_{i}$. We call all such jobs $D_{i}$.
  • If it conflicts with multiple jobs in $\mathcal{A}_{\textsc{Off}}$, then we note that at the time the job was made infeasible by $\mathcal{A}_{\textsc{Off}}$, there was a schedule $S$ that did not contain $J$ and where $J$ conflicted with a subset of $S^{\prime}\subset S$ such that $p(S^{\prime})\geq p(J)$ otherwise $J$ would be scheduled by $\mathcal{A}_{\textsc{Off}}$ instead of the jobs in $p(S^{\prime})$. As such we can divide the value of $J$ among jobs in $\mathcal{A}_{\textsc{Off}}$ such that each partial job has processing time $p(J^{\prime})\leq J_{i}$. We will refer to the partial jobs as $E_{i}$ if they are not contained within $J_{i}$ (otherwise we consider them to be in $B_{i}$). For an illustration of this process refer to figure 2.
Note that each of the conditions are in order of precedence, in other words if a job scheduled by the optimal algorithm $J_{j}$ is contained in $J_{i}$ and was also announced during $J_{i}$’s execution, then $J_{j}$ is mapped to $J_{i}$ and is part of the set $B_{i}$, not the set $C_{i}$.
Figure 1: Conflicting jobs partitioning
Figure 2: Dividing a job in conflicting with multiple jobs into partial jobs
Claim.
Let $J_{i}\in\mathcal{A}_{\textsc{Off}}$, then for any single job or partial-job $J\in A_{i}\cup D_{i}\cup E_{i}$ then $p(J)\leq p(J_{i})$
If $J\in A_{i}$ then $J=J_{i}$ and as such the inequality follows immediately.
If $J\in D_{i}$ then $J$ was announced prior to $s(J_{i})$, furthermore $J$ does not conflict with any other job scheduled by $\mathcal{A}_{\textsc{Off}}$, this means that $J$ cannot have a higher value than $J_{i}$ otherwise it would be taken instead.
If $J\in E_{i}$ then by construction this partial job must have $p(J)\leq J_{i}$
Claim.
Let $J_{i}\in\mathcal{A}_{\textsc{Off}}$, then for any single job $J\in C_{i}$ $p(J)\leq t^{-1}\cdot p(J_{i})$
This single job $J=J_{j}$ was announced during $J_{i}$’s execution and is strictly within the interval of $J_{i}$. This means the following:
$s_{i}\leq a_{j}\leq s_{i}+p_{i}$ $\text{ and }a_{j}\leq r_{j}\leq s_{i}+p_{i}$
$r_{j}-a_{j}\leq s_{i}+p_{i}-a_{j}\leq p_{i}$
$t\cdot p_{j}\leq$ $r_{j}-a_{j}\leq p_{i}$
$t\cdot p_{j}$ $\leq p_{i}$
Claim.
Let $J_{i}\in\mathcal{A}_{\textsc{Off}}$, then $p(A_{i})+p(C_{i})+p(D_{i})+p(E_{i})\leq t^{-1}\cdot p(J_{i})+p(J_{i})$
We note that none of these jobs can be contained completely with $J_{i}$. This means that at most 2 jobs from any of these sets cover the left and right endpoints of $J_{i}$’s processing interval $s_{j},s_{j}+p_{j})$. Furthermore we note that $A_{i}$ can exist if no processes the left side, otherwise $A_{i}$ could simply be scheduled by $\mathcal{A}_{\textsc{Off}}$ to get a better schedule. As such since $t^{-1}\cdot p(J_{i})\geq p(J_{i})$. $p(A_{i})+p(C_{i})+p(D_{i})+p(E_{i})\leq t^{-1}\cdot p(J_{i})+p(J_{i})$.
Claim.
Let $J_{i}\in\mathcal{A}_{\textsc{Off}}$, then $p(B_{i})\leq p(J_{i})$
The set of jobs in $B_{i}$ are all contained within the interval of $J_{i}$ and do not conflict with one another. This means that $\sum_{J_{j}\in B_{i}}\limits p_{j}\leq p_{i}$ and thus $p(B_{i})\leq p(J_{i})$.
Putting all these claims together we get that $p(A_{i})+p(C_{i})+p(D_{i})+p(E_{i})+p(B_{i})\leq p(J_{i})+t^{-1}\cdot p(J_{i})+p(J_{i})$ for every $J_{i}$. This means that $\sum_{i=0}^{n}p(A_{i})+p(C_{i})+p(D_{i})+p(E_{i})+p(B_{i})\geq\textsc{OPT}$. So $(2+t^{-1})\cdot\mathcal{A}_{\textsc{Off}}\geq\textsc{OPT}$
We have $\textsc{OPT}\cdot\frac{t}{2t+1}\leq\mathcal{A}_{\textsc{Off}}$ giving us our desired competitive ratio.
QED
Corollary 2.1.
Given sufficient $t$-advance-notice, an algorithm can reach a competitive ratio of $\frac{1}{3}$ for proportional weights real time throughput maximization.
2.2 The limitations of advance-notice
Theorem 2.2.
Suppose a sequence of jobs have $t$-advance notice, then no algorithm can achieve a better worst case asymptotic competitive ratio than $\frac{t}{2t+1}$ without preemption for proportional weight real time throughput maximization.
We will show that for any $\varepsilon>0$ and any online algorithm ALG there exists a sequence of jobs $\mathcal{J}$ such that $\textsc{OPT}\cdot\frac{t}{2t+1}+\varepsilon\leq\textsc{ALG}$.
First we fix $\gamma\leq\frac{\varepsilon(t\cdot(2t+1))}{2+t}$.
First we announce $J_{1}$ with $a_{1}=0$, $r_{1}=t$, $p_{1}=1$ and $d_{1}$ arbitrarily large.
If $J_{1}$ is never scheduled by ALG then the adversary declares it to be the last job and we have infinite competitive ratio.
If $J_{1}$ is scheduled at time $s_{1}$ then at $s_{1}$ we announce $J_{2}$ with $a_{2}=s_{1}+\gamma$, $r_{2}=s_{1}+1-2\cdot\gamma$, $p_{2}=\frac{p_{1}-\gamma}{t}$ and $d_{2}=r_{2}+p_{2}$. Note that $J_{2}$ is impossible to schedule since no preemptions is permitted. Next the adversary announces $J_{3}$ with $a_{3}=s_{1}+\gamma$, $r_{3}=s_{1}+2\cdot\gamma$ $p_{3}=\frac{\gamma}{t}$ $d_{3}=r_{3}+p_{3}$. The adversary will also announce: $J_{i}$ with $a_{i}=s_{1}+\gamma$, $r_{i}=d_{i-1}$, $p_{i}=d_{i}-r_{i}$ and $d_{i}=\min(r_{i}+\frac{\gamma}{t},r_{2})$. For $i>3$ until the interval $s_{1},s_{1}+p_{1}]$ is covered.
In this schedule, the only job that the algorithm has done is $J_{1}$ which has value $p_{1}$. Meanwhile an optimal algorithm OPT can schedule all jobs since none of them overlap except for $J_{1}$, and since $J_{1}$ has arbitrarily large slack, we can always fit it somewhere else (possibly at the beginning if the algorithm waits until the very end to schedule it).
So the value of $\textsc{ALG}=p_{1}=1$ and $\textsc{OPT}=p_{1}+p_{2}+p_{3}+\sum_{i}p_{i}=1+\frac{1-\gamma}{t}+(1-2\gamma)=\frac{2t+1-2\gamma-\gamma\cdot t}{t}=\frac{2t+1}{t}-\frac{2\cdot\gamma+\gamma\cdot t}{t}$.
So $\textsc{OPT}\cdot\frac{t}{2t+1}=\textsc{ALG}-\frac{(2\gamma+\gamma\cdot t)\cdot}{t\cdot(2t+1)}$. By our choose of $\gamma$ we get that $\textsc{OPT}\cdot\frac{t}{2t+1}+\varepsilon\leq\textsc{ALG}$
Corollary 2.2.
Regardless of how much advance-notice a sequence of jobs is guaranteed, no algorithm can ever have a competitive ratio of $\frac{1}{2}$ or better.
Corollary 2.3.
The algorithm presented in theorem 2.1 is optimal on $t\in[0,1]$
Figure 3: ALGALG vs OPTOPT schedule
3 Beyond Proportionality:
The section above allows us to understand the competitive ratio of algorithms when we are given $t$-advance-notice when weights are proportional. While these results are quite strong, restricting ourselves to proportional weights might be in some situations be too strict. We will in this section show that the performance of algorithms for cases where we are not dealing with proportional weights and show that in some of the more obvious other cases (unweighted and $C$-benevolent) regardless of $t$, an algorithm running on an set of jobs with $t$-advance-notice performs arbitrarily poorly and cannot achieve a constant competitive ratio. We will then show that if we restrict announcement times such that $a_{i}<a_{j}\implies d_{i}<d_{j}$, we can achieve $\frac{1}{2}$ competitive ratio when jobs are unweighted.
3.1 Advance-notice on C-benevolent function
From this section onwards we assume that jobs are now $J_{i}=(a_{i},r_{i},p_{i},d_{i},w_{i})\in\mathbb{R}_{\geq 0}^{5}$ our goal is now to maximize the value of $w_{i}$ for all early jobs rather than the processing time.
We borrow and adapt the following definition from Woeginger’s paper on interval scheduling [WOEGINGER19945]:
Definition 3.1.
Consider $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$. We call a sequence of jobs $f$-related if for every job $J_{i}$, $w_{i}=f(p_{i})$. We sometimes write $f(J_{i})$ as a shorthand for $f(p_{i})$
We note that proportional weights is equivalent to saying that our set of jobs is $f$-related to a linear function $f$.
Definition 3.2.
We say a function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is $C$-benevolent if the following holds:
  1. C1)
    $f(0)=0$ and $f(p)>0$ for all $p>0$
  2. C2)
    $0<\varepsilon\leq p_{1}\leq p_{2}\implies f(p_{1})+f(p_{2})\leq f(p_{1}-\varepsilon)+f(p_{2}+\varepsilon)$
  3. C3)
    $f$ is strictly increasing on $p>0$
It is clear that a linear function $f(x)=bx$ for some $b\in\mathbb{R}_{>0}$ is also $C$-benevolent. C1 follows from the fact $b\cdot x=0\iff b=0$ or $x=0$, C2 follows from additivity and C3 follows from homogeneity and $b$ being strictly positive.
In general, many of the nice properties of linear functions carry over to $C$-benevolent functions. In fact, many of the results from previous papers on interval selection and throughput apply to $C$-benevolent functions. It is why it is surprising that while we can get a finite competitive ratio with proportional weights, we cannot when the sequence of jobs is $f$-related to an arbitrary $C$-benevolent functions $f$.
Theorem 3.1.
Any algorithm ALG can have arbitrarily bad competitive ratio when the a sequence of jobs is $f$-related to a $C$-benevolent function $f$ when jobs have $t$-advance-notice for $t\in(0,1]$.
We will show that we can generate a sequence of jobs where ALG=1 and OPT=$N$ for any $N$.
The proof for this is straightforwards, we once again use an adversarial argument. Our adversary chooses $f(x)=x^{\log_{\frac{1-\varepsilon}{t}}(N)}$ to be our function $f$ for some very small epsilon. We note that this function is $C$-benevolent for $N$ sufficiently larger than $t$.
Our adversary first announces $J_{1}$ with $a_{1}=0$, $r_{1}=t$, $p_{1}=1$, $d_{1}=t+1$ and $w_{1}=f(1)$.
If our algorithm does not take the job the adversary declares it to be the last, we can assume that the algorithm takes the job.
Our adversary next announces $J_{2}$ with $a_{2}=t+\epsilon$, $r_{2}=t+1-2\varepsilon$, $p_{2}=\frac{1-\varepsilon}{t}$ $d_{2}=r_{2}+p_{2}$ which the algorithm cannot schedule.
The optimal algorithm can take $J_{2}$ which has value $f(\frac{1-\varepsilon}{t})=(\frac{1-\varepsilon}{t})^{\log_{\frac{1-\varepsilon}{t}}(N)}=N$ and the algorithm can only take $J_{1}$ which has value $f(p_{1})=f(1)=1$. We have an arbitrarily large competitive ratio as desired.
QED
3.2 Advance-notice for unweighted case
Definition 3.3.
We call a sequence of jobs “unweighted” if it is $f$-related to the function $f(x)=1$.
The unweighted case is typically one of the more commonly studied variants of real-time scheduling. We will show that for the unweighted case, we do not benefit from advanced notice.
Theorem 3.2.
Suppose a sequence of jobs is unweighted and has $t$-advanced notice. Then no algorithm $ALG$ can achieve a finite competitive ratio without preemption
Once again we show that we can generate a sequence of jobs where ALG=1 and OPT=$N$ for any $N$.
Our adversary first announces $J_{1}$ with $a_{1}=0$, $r_{1}=t$, $p_{1}=1$, $d_{1}=t+1$ and $w_{1}=f(1)$.
We then announce $N$ jobs of arbitrarily small size such that they are contained within $J_{1}$’s interval. Since all jobs have the same weight, we can make the jobs processing time small enough that the start times of all of them, along with their end times and their announcement times are all contained within $r_{1},d_{1}]$.
QED
4 Conclusion, Open Problems and Extensions
In this paper we investigated and discussed the benefits and limitations of using advance-notice in real-time online scheduling.
We first showed that there exists an algorithm which is $\frac{t}{2t+1}$-competitive on job sequences that have $t$-advance-notice for $t\in[0,1]$. While the algorithm given is of theoretical interest, it requires us to run an offline optimal algorithm on the set of known jobs whenever a new job is announced. While this is good for getting us the bounds we are looking for, since the problem of weighted job scheduling is NP-Hard [Karp1972], this algorithm is extremely computationally inefficient! As such one could wonder if there exists a more efficient algorithm for reaching this competitive ratio. Furthermore, it would be interesting to see how running our algorithm but substituting the offline algorithm for an approximation algorithm impacts the competitive ratio.
We also showed that no algorithm can achieve better than $\frac{t}{2t+1}$-competitive ratio for any $t$. This gives us a tight bound on the competitive ratio of algorithms with $t$-advance-notice for $t\in[0,1]$. But what about for $t>1$? An open question left by this paper is whether the bound holds in that case or if we can make it lower.
In the final section we showed that advance-notice does not help us much with some of the more “obvious” weight functions that people consider. These result seem to stem from a limitation of our definition of advance-notice. After all, by restricting our $(r_{i}-a_{i})$ to being proportional to $p_{i}$, it is clear that we gain information both about the processing time, but also about the weight of the job. As such, an interesting open question would be to see if we could get better results with a different definition of advance-notice that correspond more closely to the weight function used.
4.1 Acknowledgement
We are very appreciative of the feedback and encouragement from Professor Allan Borodin, and Changdao He who have provided me with valuable feedback and discussions throughout the construction of this paper.